### v3

@@ 8.2 Results
Adopted before End of Data Horizon $46\%$ $51\%$
Age at TPR (years) Age $7.71$ $(5.16)$ $9.05$ $(4.89)$
@@
Like Avery et al. (2009), we model children’s time to placement using a Cox proportional hazards model (Cox 1972). The Cox proportional hazards model includes a baseline hazards function for how the likelihood of a placement changes over time and a parameter for each covariate that affects the baseline hazard. The following characteristics were available in both AFCARS and the platform data, allowing us to control for them in the hazards model (for discrete variables, one of the categories is omitted for statistical identification):
@@
    Black or African-American (versus neither Black nor African-American, the omitted category);
@@
    federal fiscal year (e.g., October 1, 2014, to September 30, 2015, for FY2015) of the TPR order, with the fiscal year from October 1, 2020 to September 30, 2021 or FY2021 as the omitted category.
Table 1 provides summary statistics for these variables. Because the most recent AFCARS files are only available through September 30, 2021, we additionally excluded 90 cases that began on the platform after that date due to our inability to estimate their entry-year effect from the AFCARS data. The platform’s population of children tends to be older, more male, and more likely to have a clinically diagnosed disability — all factors associated with greater difficulty in placing children. A lower percentage of children on the platform tended to be white, although discrepancies exist in how the AFCARS data and the platform categorize race. To be conservative in our estimation of the platform’s performance, we accommodate race through a variable for children identified as Black; i.e., counting some of the children served by the platform in the “Other” category as part of a racial group associated in other research with lower rates of permanency would lower the performance benchmark for the platform.
Table 2 shows how basic demographic characteristics affect the time to placement using the AFCARS dataset for Florida children. One model includes controls for the fiscal year in which TPR occurred, and a second simpler model only includes demographic characteristics. Due to its heightened predictive power, we choose to use Model 1 for predicting child outcome probabilities and establishing the AFCARS benchmark. Controlling for the fiscal year of the start of the search, covariates associated with a faster placement are being female and having Hispanic ethnicity; factors associated with a slower time to or diminished likelihood of adoption are being older, having a disability, and being Black. Using this model, we collected the necessary data on each child served by the platform, and we predicted the likelihood of a finalized adoption at monthly intervals up to the duration for which the child was in need of an adoption placement. Children listed before the COVID years, FY2020 and the control year FY2021, were likely to be adopted faster than those with TPR in FY2020 and FY2021. If a child was adopted, the predictions continued assuming the child would have been active until February 1, 2023, or the child turned 18, whichever came first.
Table 2: Cox Proportional Hazards Model of Adoptions
@@
Black $0.768^{\ast\ast\ast}$ $0.772^{\ast\ast\ast}$
$(-8.39)$ $(-8.21)$
Hispanic $0.827^{\ast\ast\ast}$ $0.833^{\ast\ast\ast}$
$(-4.31)$ $(-4.14)$
Age at TPR (years) $0.889^{\ast\ast\ast}$ $0.894^{\ast\ast\ast}$
$(-9.59)$ $(-9.15)$
@@
TPR in FY2016 $1.959^{\ast\ast\ast}$
$(8.45)$
TPR in FY2017 $2.131^{\ast\ast\ast}$
$(9.63)$
TPR in FY2018 $1.881^{\ast\ast\ast}$
$(8.18)$
TPR in FY2019 $1.526^{\ast\ast\ast}$
$(5.43)$
TPR in FY2020 $1.046$
$(0.55)$
N 10,284 10,284
Concordance $0.680$ $0.665$
Log-likelihood ratio test 1965.47 on 12 d.f. 1683.39 on 6 d.f.
@@
By aggregating the predicted adoption probabilities of all children served by the platform, we can establish a benchmark against which to compare the actual adoptions of children on the platform. Let $\mathcal{C}$ represent the set of children. For any child $i$ with attributes $X_{i}$, $i\in\mathcal{C}$, we can predict the survival probability using the adoption probability function $\hat{\pi}(X_{i},t)$ that child $i$ has a finalized adoption within $t$ years. The child’s time until adoption finalization, unsuccessful search conclusion, or February 1, 2023, is recorded as $\tau_{i}$, and $\tau^{e}_{i}$ extends $\tau_{i}$ to include an extended search horizon until the child turns 18 for children who are adopted. We also have a delay value $\tau_{i}^{d}$ for the time between TPR and registration on the platform. This time duration is estimated as the platform allows caseworkers to register children and select time buckets — for which we use the mid-point — from the time since TPR instead of a specific date. We provide an alternate analysis in which we assume $\tau_{i}^{d}=0$ in online Appendix F. Thus, we calculate a conditional survival probability for child $i$ as
$\tilde{\pi}(X_{i},t,\tau_{i}^{d}):=\frac{\hat{\pi}(X_{i},t)-\hat{\pi}(X_{i},\tau_{i}^{d})}{1-\hat{\pi}(X_{i},\tau_{i}^{d})},$ (13)
@@
which provides the AFCARS benchmark $\mu(t)$ for the number of expected matches by time $t$:
$\mu(t):=\sum_{i\in\mathcal{C}}\tilde{\pi}(X_{i},\min\{t,\tau^{e}_{i}\},\tau_{i}^{d}).$ (14)
@@
Figure 5 shows the adoptions achieved through the platform compared to this benchmark from the AFCARS proportional hazards model. The results show that the platform has outperformed the commonly used two-year and three-year search window benchmarks. For children listed on the platform before October 1, 2021, 138 adoptions were finalized within two years — 123 enabled by the platform and 15 through other channels — while the predicted number for children on the platform was only 124.5. At the three-year mark, this difference between adoptions achieved and the predicted number extends to an extra 24.3 adopted children, or a 17% increase over the benchmark. However, the platform’s performance fell short of the AFCARS benchmark one year after the search began, even though the platform not only made up for this gap but surpassed the benchmark at the two-year and three-year marks. The difference in performance may be attributed to a limitation of the AFCARS datasets mentioned above: some cases included in the benchmark might not have actually required a search for a family but rather had a potentially faster and easier non-relative adoptive placement. Despite this disadvantage, the platform’s over-performance over longer time horizons may indicate that the technology helps caseworkers to be more persistent in their search efforts for hard-to-place children. When Avery (2000) investigated the longest-waiting children in New York, she found case workers to be pessimistic about the children’s chances at adoption and that the case workers failed to use search tools available to them. However, this case study shows that providing leads of potential adoptive families may encourage caseworkers to persist in finding families.

### v4

@@ 9.1 Statistical Approaches
Adopted before End of Data Horizon $46\%$ $59\%^{\dagger}$
Age at TPR (years) Age $7.71$ $(5.16)$ $8.39$ $(4.73)$
@@
Includes 144 adoptions via the platform and 21 children listed on the platform but adopted by
non-relative/non-foster families found off-platform.
Both of our statistical approaches model children’s time to placement using a Cox proportional hazards model (Cox 1972). The Cox proportional hazards model includes a baseline hazard function that describes how the likelihood of a placement changes over time, as well as a parameter for each covariate that affects the baseline hazard. The following characteristics were available in both AFCARS and the platform data, allowing us to control for them in the hazards model (for discrete variables, one of the categories is omitted for statistical identification):
@@
    Black or African-American (versus no such designation, the omitted category);
@@
    federal fiscal year (e.g., October 1, 2014, to September 30, 2015, for FY2015) of the termination of the parental rights (TPR) order, with FY2015 as the omitted category.
Focusing on Model 1 — the Cox proportional hazards model without a platform effect that is used for benchmarking — let $t=0$ denote the TPR date for some child $i$. The hazard of adoption finalization at time $t$ is given by the semi-parametric Cox model
${\small h\!\left(t\mid X_{i}\right)=h_{0}(t)\;\exp\!\Bigl{(}}$ ${\small\beta_{1}\text{Female}_{i}+\beta_{2}\text{Black}_{i}+\beta_{3}\text{Hispanic}_{i}+\beta_{4}\text{Disability}_{i}+\beta_{5}\text{Age}_{i}+\beta_{6}\text{Age}_{i}^{2}}$
$\textstyle{\small+\!\!\sum_{y=2016}^{2021}\!\beta_{y}\,\mathbbm{1}\{\text{FY}_{i}=y\}\Bigr{)},}$ (13)
where $h_{0}(t)$ is a baseline hazard rate common to all children [^11] The baseline hazard rate is estimated using the methodology of Breslow (1972). and $X_{i}$ collects the covariates listed above. The indicator $\mathbbm{1}\{\text{FY}_{i}=y\}$ flags the fiscal year in which TPR occurred. Estimation relies on Cox’s partial likelihood, which is obtained using the hazard rates of all children. Maximizing the log-partial likelihood yields the coefficient estimates. An exponentiated coefficient $\exp(\beta_{k})>1$ indicates that covariate $k$ is associated with a faster‐than‐baseline adoption rate, whereas $\exp(\beta_{k})<1$ signals a slower rate. We explain how the estimated model in Equation (13) changes when the platform effect is included for the second approach in Section 9.3.
9.2 Benchmark Against Statewide Outcomes
Using only the statewide AFCARS data, we construct a benchmark for the outcomes of children served by the platform. Model 1 of Table 2 shows how demographic factors affect the time to placement using the AFCARS dataset for Florida children, i.e., the maximum-likelihood estimation of the Cox model in Equation (13). Controlling for the fiscal year in which the search started, covariates associated with a faster placement are being female and having Hispanic ethnicity; factors associated with a slower time to or diminished likelihood of adoption are being older, having a disability, and being Black. Using this fitted model, we collected the necessary covariate data on each child served by the platform and predicted the likelihood of a finalized adoption at monthly intervals, as explained below. If a child is adopted, the benchmark adoption probability continues accumulating as if the child’s search continued until the earlier date of February 1, 2023, or the child’s 18th birthday.
Table 2: Cox Proportional Hazards Models of Time Until Adoption
@@ 9.2 Benchmark Against Statewide Outcomes
Black $0.768^{\ast\ast\ast}$ $0.769^{\ast\ast\ast}$
$(-8.379)$ $(-8.441)$
Hispanic $0.827^{\ast\ast\ast}$ $0.841^{\ast\ast\ast}$
$(-4.313)$ $(-3.963)$
Age at TPR (years) $0.889^{\ast\ast\ast}$ $0.891^{\ast\ast\ast}$
$(-9.563)$ $(-9.507)$
@@
Platform Cases 0 279
N 10,286 10,565
Concordance $0.680$
Log-likelihood ratio test 1969.148 2038.169
on 12 d.f. on 13 d.f.
@@
By adding the predicted adoption probabilities of all children served by the platform, we can establish a benchmark against which to compare the actual adoptions of children on the platform. Let $\mathcal{C}$ represent the set of children on the platform. For any child $i\in\mathcal{C}$ with attributes $X_{i}$, we can use the survival probability to predict the adoption probability function using the estimated hazard function $\hat{h}(\cdot|X_{i})$ in Model 1
$\hat{\pi}(X_{i},t):=1-\exp\!\left(-\int_{0}^{t}\hat{h}\!(u\mid X_{i})\,\mathrm{d}u\right)$ (14)
for the likelihood that child $i$ has a finalized adoption within $t$ years.
A child’s maximum possible search horizon from TPR until the child turns 18 is denoted by $\tau_{i}^{e}$. We use $\tau_{i}^{d}$ to represent the time between the TPR date and the earliest possible adoption finalization date after registration on the platform. We calculate $\tau_{i}^{d}$ as the time between the case creation date and the TPR date, plus an additional three months to account for the legally required period in Florida during which a child must reside with the adoptive family before the adoption is finalized. [^12] We provide an alternate analysis in which we assume $\tau_{i}^{d}=0$, as well as an analysis excluding platform children whose placement was identified through other channels in online Appendix I.2.
To account for the time that a child might have already been eligible for adoption before being listed on the platform, we calculate a conditional survival probability for child $i$ at any time $t>\tau_{i}^{d}$ since TPR as
$\tilde{\pi}(X_{i},t,\tau_{i}^{e},\tau_{i}^{d}):=\frac{\hat{\pi}(X_{i},\min\{t,\tau^{e}_{i}\})-\hat{\pi}(X_{i},\tau_{i}^{d})}{1-\hat{\pi}(X_{i},\tau_{i}^{d})},$ (15)
@@
which provides the AFCARS benchmark $\hat{\mu}(t)$ for the number of expected matches by time $t$:
$\hat{\mu}(t):=\textstyle\sum_{i\in\mathcal{C}}\tilde{\pi}(X_{i},t,\tau_{i}^{e},\tau_{i}^{d}).$ (16)
@@
Figure 5 shows the adoptions achieved through the platform compared to this benchmark from the AFCARS proportional hazards model. The results show that the platform has outperformed the commonly used two-year and three-year search window benchmarks. For children listed on the platform before October 1, 2021, 138 adoptions were finalized within two years — 123 enabled by the platform and 15 through other channels — while the predicted number for children on the platform was only 117.2. At the three-year mark, this difference between adoptions achieved and the predicted number extends to an extra 31.6 adopted children, or a 24% increase over the benchmark. However, the platform’s performance fell short of the AFCARS benchmark one year after the search began, even though the platform not only made up for this gap but surpassed the benchmark at the two-year and three-year marks. The difference in performance may be attributed to a limitation of the AFCARS datasets mentioned above: some cases included in the benchmark might not have actually required a search for a family but rather had a potentially faster and easier non-relative adoptive placement. One subject-matter expert shared that it is common for foster families to network and recruit on behalf of children in their care who might need adoptive placements; these children might have an adoptive resource identified at TPR who is not a relative or foster parent and be included in our benchmark dataset.
Despite this disadvantage, the platform’s over-performance over longer time horizons may indicate that the technology helps caseworkers to be more persistent in their search efforts for hard-to-place children. When Avery (2000) investigated the longest-waiting children in New York, caseworkers were found to be pessimistic about the children’s chances at adoption. They also neglected to use the search tools available to them. However, in addition to the evidence of usage from the actual placements generated by the platform, a survey of the platform’s 73 active users in Florida suggests a positive assessment of caseworker satisfaction with the new CS-driven system. A total of 51 respondents — mostly caseworkers but also some supervisors and recruitment specialists — responded to the survey in 2023. About 30% of the respondents had been in their current role for at least five years, so many of the respondents were familiar with other search methods besides the platform. When asked how satisfied they were with the platform, approximately 33% were very satisfied, and 31% were somewhat satisfied. Only 5% were very dissatisfied, and only 9% were somewhat dissatisfied. Thus, this case study shows that providing leads of potential adoptive families may encourage caseworkers to persist in finding families.